Evaluate M- )dx, using two points Gauss Lengendre Quadrature.
Answer #1

By Two-Point Gauss-Legendre Rule, if f is continuous on [-1, 1] then

\int_{-1}^{1} f(x) dx \approx G_2(f) = f\left ( -\frac{1}{\sqrt{3}} \right ) + f\left ( \frac{1}{\sqrt{3}} \right )

\int_{-1}^{1} \sqrt{1-x^2} dx \approx G_2(f) = \sqrt{1-\left ( \frac{-1}{\sqrt{3}} \right )^2} + \sqrt{1-\left ( \frac{1}{\sqrt{3}} \right )^2}

= \sqrt{\frac{2}{3}} +\sqrt{\frac{2}{3}} = 2\sqrt{\frac{2}{3}} = 1.632993